OpenStudy (anonymous):

Find f'(x) when f(x)=x^2+x , using the definition of the derivative, if that makes sense..

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (angela210793):

F'(x)= 2x+1

OpenStudy (angela210793):

(x^n)'=X^(n-1)...u know this?

OpenStudy (anonymous):

Kinda :/

OpenStudy (anonymous):

x^n=nx^(n-1)

OpenStudy (angela210793):

yea i 4got tht...sorry ur right

OpenStudy (anonymous):

no problem...

OpenStudy (angela210793):

^_^

OpenStudy (anonymous):

it says to use the definition, so i think you are supposed to write \[lim_{h->0}\frac{f(x+h)-f(x)}{h}=lim_{h->0}\frac{(x+h)^2+(x+h)-(x^2+x)}{h}\]

OpenStudy (anonymous):

if you just need the answer then angela gave it to you but if you have to write it to hand in you need to use the definition, not the power rule.

OpenStudy (anonymous):

\[\frac{x^2+2xh+h^2+x+h-x^2-x}{h}=\frac{2xh+h+h^2}{h}=2x+1+h\]and now let h to to zero to get your answer.

OpenStudy (anonymous):

Thanks! :)